### v1

@@ Abstract
While large language models (LLMs) achieve near-perfect scores on medical licensing exams, these evaluations inadequately reflect the complexity and diversity of real-world clinical practice. We introduce MedHELM, an extensible evaluation framework for assessing LLM performance for medical tasks with three key contributions. First, a clinician-validated taxonomy spanning 5 categories, 22 subcategories, and 121 tasks developed with 29 clinicians. Second, a comprehensive benchmark suite comprising 35 benchmarks (17 existing, 18 newly formulated) providing complete coverage of all categories and subcategories in the taxonomy. Third, a systematic comparison of LLMs with improved evaluation methods (using an LLM-jury) and a cost-performance analysis. Evaluation of 9 frontier LLMs, using the 35 benchmarks, revealed significant performance variation. Advanced reasoning models (DeepSeek R1: 66% win-rate; o3-mini: 64% win-rate) demonstrated superior performance, though Claude 3.5 Sonnet achieved comparable results at 40% lower estimated computational cost. On a normalized accuracy scale (0-1), most models performed strongly in Clinical Note Generation (0.73-0.85) and Patient Communication & Education (0.78-0.83), moderately in Medical Research Assistance (0.65-0.75), and generally lower in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Our LLM-jury evaluation method achieved good agreement with clinician ratings (ICC = 0.47), surpassing both average clinician–clinician agreement (ICC = 0.43) and automated baselines including ROUGE-L (0.36) and BERTScore-F1 (0.44). Claude 3.5 Sonnet achieved comparable performance to top models at lower estimated cost. These findings highlight the importance of real-world, task-specific evaluation for medical use of LLMs and provides an open source framework to enable this.
@@ 1 Introduction
However, there is a large gap between performance on medical knowledge benchmarks and readiness for real-world deployment due to three key limitations in these existing benchmarks: (1) Questions do not match real-world settings – Existing benchmarks rely on synthetic vignettes or narrowly-scoped exam questions, failing to capture key aspects of real diagnostic processes such as extracting relevant details from patient records [5, 6]. (2) Limited use of real-world data – Only 5% of LLM evaluations use real-world electronic health record (EHR) data [7]. EHRs contain ambiguities, inconsistencies, and domain-specific shorthand that synthetic data cannot replicate. (3) Limited task diversity – Around 64% of LLM evaluations in healthcare focus only on medical licensing exams and diagnostic tasks [7], ignoring essential hospital operations such as administrative tasks (e.g., generating prior authorization letters, identifying billing codes), clinical documentation (e.g., writing progress notes or discharge instructions), and patient communication (e.g., asynchronous messaging through electronic patient portals) [8].
@@
To address these limitations, we introduce MedHELM (Holistic Evaluation of Large Language Models for Medical Tasks), an extensible evaluation framework for assessing LLM performance in completing medical tasks (Figure 1). Inspired by the HELM project’s standardized cross-domain evaluations [10], using MedHELM  we evaluate 9 LLMs using 35 distinct benchmarks covering all 22 subcategories of medical tasks, focusing on clinicians’ day-to-day activities beyond just taking licensing exams. We assess performance using benchmark-appropriate metrics (i.e., exact match for closed-ended benchmarks, LLM-jury where three LLMs evaluate responses using tailored rubrics for open-ended benchmarks with demonstrated agreement to clinician ratings) as well as estimated computational cost to provide practical deployment insights. Our primary contributions are:
@@
    Clinician-validated taxonomy: A 5-category, 22-subcategory, 121-task taxonomy developed with 29 clinicians. Clinicians achieved a 96.7% agreement rate when mapping subcategories to appropriate categories, validating the clear and discrete nature of the taxonomy. We present the complete taxonomy in the Results and Appendix sections.
@@
    A benchmark suite with full taxonomy coverage: A collection of 35 benchmarks spanning all 22 subcategories of medical tasks. This includes 17 existing benchmarks, 5 re-formulated benchmarks based on existing datasets, and 13 new benchmarks. [^1] For privacy and regulatory compliance, as well as to prevent inclusion in LLM training data, 14 datasets are not publicly released.
@@
    Comparative evaluation of models along with cost-performance analysis: A systematic evaluation shows that reasoning models achieve the highest overall performance. For example, DeepSeek R1 and o3-mini score the best in Clinical Note Generation (0.85) while o3-mini leads in Clinical Decision Support (0.72) and Medical Research Assistance (0.75)  [^2] Note: Values in parentheses represent scores on a normalized accuracy scale (0-1).. On a subset of samples from two open-ended benchmarks (MEDIQA-QA & ACI-Bench), our LLM-jury approach achieves 0.47 agreement (as measured by intraclass correlation coefficient) with clinicians, beating average clinician-to-clinician agreement (0.43) and outperforming ROUGE-L (text overlap metric, 0.36) and BERTScore-F1 (semantic similarity metric, 0.44). A cost-performance analysis identifies Claude 3.5 Sonnet as providing comparable results with lower estimated computational cost.
The MedHELM framework addresses a critical need in the testing and evaluation of AI for medical use by providing consistent, real-world evaluation standards for medical (and healthcare) applications of LLMs. This framework benefits three key stakeholder groups: (1) Healthcare systems evaluating LLMs for certain tasks, (2) AI developers identifying performance gaps across medical tasks, and (3) Researchers developing methods to reproducibly measure LLM capabilities on medical tasks. To foster collaborative improvement of such AI evaluation, we provide an openly accessible leaderboard with current benchmarking results and share the codebase  [^3] https://github.com/stanford-crfm/helm, with documentation  [^4] https://crfm-helm.readthedocs.io/en/latest/medhelm/ for contributing new datasets, evaluation metrics, and benchmarking custom models. By standardizing terminology and evaluation methods across the task taxonomy, MedHELM establishes a foundation for reproducible and real-world assessment of AI capabilities in medicine.
Figure 1: This figure illustrates: (a) a clinician-validated taxonomy organizing 121 medical tasks into 5 categories and 22 subcategories; (b) a suite of benchmarks that map existing benchmarks to this taxonomy and introduces new benchmarks for complete coverage; and (c) an evaluation comparing reasoning and non-reasoning LLMs, with model rankings, LLM jury based evaluation of open-ended benchmarks, and cost-performance analysis
@@ 2.1 Clinician Validation of the Taxonomy
In a structured review process, clinicians evaluated the initial taxonomy comprising 5 categories, 21 subcategories, and 98 tasks. When asked to assign each subcategory to its appropriate top-level category, clinicians correctly matched 96.7% of subcategories to their intended categories. The clinicians rated the comprehensiveness of the proposed tasks at a mean of 4.21/5 (n = 29) and provided 107 comments with suggestions for improvement. Based on this feedback, we refined task definitions and expanded the taxonomy to its final form: 5 categories, 22 subcategories, and 121 tasks. An overview of the final taxonomy is shown in Figure 2, with the complete list provided in Appendix A.
Figure 2: Overview of the final taxonomy comprising 5 main categories and 22 subcategories.
@@ 2.2 Overview of the Benchmark Suite
Our 35 benchmarks span all 22 subcategories, providing full coverage over categories and subcategories in our taxonomy (Table 7). The benchmark suite comprises 17 existing benchmarks, 5 re-formulated benchmarks derived from previously unevaluated medical datasets, and 13 new benchmarks, out of which 12 are EHR-based. The suite includes 13 open-ended benchmarks (requiring free-text generation) and 22 closed-ended benchmarks (with predefined answer choices). Access levels are designated as 14 public, 7 gated (i.e. requiring approval) and 14 private.
Clinical Decision Support is the most represented category with 10 benchmarks, followed by Patient Communication (8), Clinical Note Generation and Medical Research Assistance (6), and Administration & Workflow (5). The distribution of benchmarks across subcategories is uneven, with 15 subcategories containing a single benchmark and the remaining 7 subcategories containing between 2 and 5 benchmarks each.
@@ 2.3.1 Overall Performance
Table 1 compares models using win-rate and macro-average performance metrics (defined in the table caption). DeepSeek R1 performed best, winning 66% of head-to-head comparisons with a macro-average of 0.76 and low win standard deviation (0.10). o3-mini followed with a 64% win rate and the highest macro-average (0.78), driven by strong performance in benchmarks in the clinical decision support category.
@@
DeepSeek R1 0.66 0.10 0.76 0.22
o3-mini (2025-01-31) 0.64 0.16 0.78 0.19
Claude 3.7 Sonnet (20250219) 0.64 0.13 0.74 0.21
Claude 3.5 Sonnet (20241022) 0.63 0.14 0.74 0.21
GPT-4o (2024-05-13) 0.57 0.17 0.75 0.19
Gemini 2.0 Flash 0.42 0.17 0.72 0.21
GPT-4o mini (2024-07-18) 0.39 0.18 0.73 0.20
Llama 3.3 Instruct (70B) 0.30 0.13 0.71 0.22
Gemini 1.5 Pro (001) 0.24 0.08 0.69 0.21
@@
We present every model’s normalized score from each of the 35 benchmarks as a heatmap in Figure 3, where darker green indicates higher performance. We see that models perform worse on benchmarks such as MedCalc-Bench (calculating medical values from patient notes), EHRSQL (generating SQL queries from natural language instructions for clinical research—originally intended as code generation dataset), and MIMIC-IV Billing Code (assigning ICD-10 codes to clinical notes). The best performance is seen for the NoteExtract benchmark (extracting specific information from clinical notes). Broader category-level trends are described below.
Figure 3: Heatmap of normalized scores (0–1) for each model (rows) across 35 benchmarks (columns). Dark green indicates high performance; dark red indicates low performance. The statistical significance of model differences varies by benchmark, with an analysis of minimum detectable effects shown in Appendix E. Metrics include EM as exact match, Jury Score as the average normalized score from 3 frontier LLMs, MedFlagAcc as binary accuracy for detecting presence of errors, MedCalcAcc as exact match or thresholded match depending on question type, and EHRSQLExecAcc as execution accuracy of generated code against a target output.
@@
Figure 4 presents performance scores by the 5 top-level categories in the taxonomy. Most models achieve their highest mean scores in Clinical Note Generation (0.73-0.85, with Gemini 1.5 Pro as an outlier at 0.56) and Patient Communication & Education (0.78-0.83), moderate performance in Medical Research Assistance (0.65-0.75), and generally lower scores in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Performance variations across categories reflect distinct task-based challenges: free-text generation tasks (Clinical Note Generation, Patient Communication) leverage models’ natural language strengths, while structured reasoning tasks (Clinical Decision Support, Administration) require domain-specific knowledge integration and logical inference. These patterns have important implications for selective deployment strategies in healthcare settings.
At the model level, DeepSeek R1 and o3-mini lead across most categories, with DeepSeek R1 excelling in Clinical Note Generation (0.85, tied with o3-mini) and Patient Communication & Education (0.82), while o3-mini leads in Clinical Decision Support (0.72) and Medical Research Assistance (0.75). GPT-4o achieves the highest score in Patient Communication & Education (0.83). The Claude Sonnet series (3.5 and 3.7) shows consistent performance across categories, with scores of 0.78-0.81 in Clinical Note Generation and Patient Communication & Education. The Gemini series of models shows significant performance variability across categories, performing well in Patient Communication & Education tasks (0.80) but poorly in other categories. This likely reflects stricter safety guardrails leading to more question refusals. The open-source Llama 3.3 performs well in Patient Communication & Education (0.82) but shows the lowest score in Administration & Workflow (0.53), indicating areas for future improvement.
Figure 4: Mean normalized scores (0-1 scale) across the 5 categories for all evaluated models. Darker green represents higher scores. Models are ordered by mean win rate from top (highest) to bottom (lowest), while categories are arranged left to right.
@@ 2.3.3 Cost–Performance Analysis
We estimated the cost (USD) of evaluating each model based on publicly listed pricing as of 05/12/2025, using the total input tokens and maximum output tokens consumed during benchmark runs and LLM-jury evaluation. These costs are an upper-bound estimate, since models may generate fewer tokens than the maximum allowed output. We plotted the mean win-rate against the cost (Figure 5, with a detailed breakdown in Table 3, and inference costs in Appendix G). As expected, non-reasoning models—GPT-4o mini (\$805) and Gemini 2.0 Flash (\$815)—incurred the lowest costs and achieved win-rates of 0.39 and 0.42, respectively. Open-source Llama 3.3 Instruct (\$940) had a 0.30 win-rate, while Gemini 1.5 Pro (\$1,130) reached 0.24. Reasoning models incurred higher costs, DeepSeek R1 (\$1,806) and o3-mini (\$1,722), with win-rates of 0.66 and 0.64 respectively. Claude 3.5 Sonnet (\$1,571) and Claude 3.7 Sonnet (\$1,537) provide a good cost-performance balance, achieving $\sim 0.63$ win-rate at reduced costs.
Figure 5: Scatter plot of mean win-rate (y-axis) versus estimated computational cost (x-axis) for each of the 9 models across 35 benchmarks. Each point represents a model, with the position indicating the relationship between performance (y-axis) and total cost of evaluation including benchmark runs and evaluation by LLM-jury (x-axis). Costs represent upper-bound estimates based on maximum output token usage.
@@
Table 3: Comparison of LLMs by architecture, access type, and token usage metrics across MedHELM. Benchmark tokens denote total input/output tokens in completing the medical task represented by the benchmark; jury tokens are tokens used for open-ended evaluations via the LLM-jury. Total cost reflects the estimated per-model expense of running a MedHELM evaluation across 35 benchmarks and represents an upper bound based on maximum output token usage. For 9 models and 35 benchmarks in the current MedHELM suite, one update to the leaderboard is estimated to cost $11,739.07$.

### v2

@@ Abstract
While large language models (LLMs) achieve near-perfect scores on medical licensing exams, these evaluations inadequately reflect the complexity and diversity of real-world clinical practice. We introduce MedHELM, an extensible evaluation framework for assessing LLM performance for medical tasks with three key contributions. First, we present a clinician-validated taxonomy spanning five categories, 22 subcategories, and 121 tasks developed with 29 clinicians. Second, we develop a comprehensive benchmark suite comprising 35 benchmarks (17 existing, 18 newly formulated) providing complete coverage of all categories and subcategories in the taxonomy. Third, we conduct a systematic comparison of LLMs with improved evaluation methods (using an LLM-jury) and a cost-performance analysis. Evaluation of nine frontier LLMs, using the 35 benchmarks, revealed significant performance variation. Advanced reasoning models (DeepSeek R1: 66% win-rate; o3-mini: 64% win-rate) demonstrated superior performance, though Claude 3.5 Sonnet achieved comparable results at 40% lower estimated computational cost. On a normalized accuracy scale (0-1), most models performed strongly in Clinical Note Generation (0.74-0.85)  and Patient Communication & Education (0.76-0.89), moderately in Medical Research Assistance (0.65-0.75) and Clinical Decision Support (0.61-0.76), and lower in Administration & Workflow (0.53-0.63). Our LLM-jury evaluation method achieved good agreement with clinician ratings (ICC = 0.47), surpassing both average clinician–clinician agreement (ICC = 0.43) and automated baselines, including ROUGE-L (0.36) and BERTScore-F1 (0.44). Claude 3.5 Sonnet achieved comparable performance to top models at lower estimated cost. These findings highlight the importance of real-world, task-specific evaluation for medical use of LLMs and provides an open source framework to enable this.
@@ 1 Introduction
However, there is a large gap between performance on medical knowledge benchmarks and readiness for real-world deployment due to three key limitations in these existing benchmarks [5]: (1) Questions do not match real-world settings – Existing benchmarks rely on synthetic vignettes or narrowly-scoped exam questions, failing to capture key aspects of real diagnostic processes such as extracting relevant details from patient records [6, 7]. (2) Limited use of real-world data – Only 5% of LLM evaluations use real-world electronic health record (EHR) data [8]. EHRs contain ambiguities, inconsistencies, and domain-specific shorthand that synthetic data cannot replicate. (3) Limited task diversity – Around 64% of LLM evaluations in healthcare focus only on medical licensing exams and diagnostic tasks [8], ignoring essential hospital operations such as administrative tasks (e.g., generating prior authorization letters, identifying billing codes), clinical documentation (e.g., writing progress notes or discharge instructions), and patient communication (e.g., asynchronous messaging through electronic patient portals) [9].
@@
To address these limitations, we introduce MedHELM (Holistic Evaluation of Large Language Models for Medical Tasks), an extensible evaluation framework for assessing LLM performance in completing medical tasks (Figure 1). Inspired by the HELM project’s standardized cross-domain evaluations [11], using MedHELM  we evaluate nine LLMs using 35 distinct benchmarks covering all 22 subcategories of medical tasks, focusing on clinicians’ day-to-day activities beyond just taking licensing exams. We assess performance using benchmark-appropriate metrics (i.e., exact match for closed-ended benchmarks, LLM-jury where three LLMs evaluate responses using tailored rubrics for open-ended benchmarks with demonstrated agreement to clinician ratings) as well as estimated computational cost to provide practical deployment insights. Our primary contributions are:
@@
    Clinician-validated taxonomy: A five-category, 22-subcategory, 121-task taxonomy developed with 29 clinicians. Clinicians achieved a 96.7% agreement rate when mapping subcategories to appropriate categories, validating the clear and discrete nature of the taxonomy. We present the complete taxonomy in the Results and Appendix sections.
@@
    A benchmark suite with full taxonomy coverage: A collection of 35 benchmarks spanning all 22 subcategories of medical tasks. This includes 17 existing benchmarks, five re-formulated benchmarks based on existing datasets, and 13 new benchmarks. [^1] For privacy and regulatory compliance, as well as to prevent inclusion in LLM training data, 14 datasets are not publicly released.
@@
    Comparative evaluation of models along with cost-performance analysis: A systematic evaluation shows that reasoning models achieve the highest overall performance.
The MedHELM framework addresses a critical need in the testing and evaluation of AI for medical use by providing consistent, real-world evaluation standards for medical (and healthcare) applications of LLMs. This framework benefits three key stakeholder groups: (1) Healthcare systems evaluating LLMs for specific tasks, (2) AI developers identifying performance gaps across medical tasks, and (3) Researchers developing methods to reproducibly measure LLM capabilities on medical tasks. To foster collaborative improvement of such AI evaluation, we provide an openly accessible leaderboard  [^2] https://crfm.stanford.edu/helm/medhelm/v2.0.0/ with current benchmarking results and share the codebase  [^3] https://github.com/stanford-crfm/helm, with documentation  [^4] https://crfm-helm.readthedocs.io/en/latest/medhelm/ for contributing new datasets, evaluation metrics, and benchmarking custom models. By standardizing terminology and evaluation methods across the task taxonomy, MedHELM establishes a foundation for reproducible and real-world assessment of AI capabilities in medicine.
Figure 1: This figure illustrates: (a) a clinician-validated taxonomy organizing 121 medical tasks into five categories and 22 subcategories; (b) a suite of benchmarks that map existing benchmarks to this taxonomy and introduces new benchmarks for complete coverage; and (c) an evaluation comparing reasoning and non-reasoning LLMs, with model rankings, LLM jury based evaluation of open-ended benchmarks, and cost-performance analysis
@@ 2.1 Clinician Validation of the Taxonomy
In a structured review process, 29 clinicians evaluated the initial taxonomy comprising five categories, 21 subcategories, and 98 tasks. When asked to assign each subcategory to its appropriate top-level category, clinicians correctly matched 96.7% of subcategories to their intended categories. The clinicians rated the comprehensiveness of the proposed tasks at a mean of 4.21/5 (n = 29) and provided 107 comments with suggestions for improvement. Based on this feedback, we refined task definitions and expanded the taxonomy to its final form: five categories, 22 subcategories, and 121 tasks. An overview of the final taxonomy is shown in Figure 2, with the complete list provided in Appendix A.
Figure 2: Overview of the final taxonomy comprising five main categories and 22 subcategories.
@@ 2.2 Overview of the Benchmark Suite
Our 35 benchmarks span all 22 subcategories, providing full coverage over categories and subcategories in our taxonomy (Table 7). The benchmark suite comprises 17 existing benchmarks, five re-formulated benchmarks derived from previously unevaluated medical datasets, and 13 new benchmarks, out of which 12 are EHR-based. The suite includes 13 open-ended benchmarks (requiring free-text generation) and 22 closed-ended benchmarks (with predefined answer choices). Access levels are designated as 14 public, seven gated (i.e. requiring approval), and 14 private.
Clinical Decision Support is the most represented category with ten benchmarks, followed by Patient Communication (eight), Clinical Note Generation and Medical Research Assistance (six), and Administration & Workflow (five). The distribution of benchmarks across subcategories is uneven, with 15 subcategories containing a single benchmark and the remaining seven subcategories containing between two and five benchmarks each.
@@ 2.3.1 Overall Performance
Table 1 compares models using win-rate and macro-average performance metrics (defined in the table caption). DeepSeek R1 performed best, winning 66% of head-to-head comparisons with a macro-average of 0.75 and low win standard deviation (0.10). o3-mini followed with a 64% win rate and the highest macro-average (0.77), driven by strong performance in benchmarks in the clinical decision support category.
@@
DeepSeek R1 0.66 0.10 0.75 0.22
o3-mini (2025-01-31) 0.64 0.16 0.77 0.18
Claude 3.7 Sonnet (20250219) 0.64 0.13 0.73 0.21
Claude 3.5 Sonnet (20241022) 0.63 0.14 0.73 0.21
GPT-4o (2024-05-13) 0.57 0.17 0.73 0.18
Gemini 2.0 Flash 0.42 0.17 0.70 0.21
GPT-4o mini (2024-07-18) 0.39 0.18 0.71 0.20
Llama 3.3 Instruct (70B) 0.30 0.13 0.69 0.22
Gemini 1.5 Pro (001) 0.24 0.08 0.67 0.21
@@
We present every model’s normalized score from each of the 35 benchmarks as a heatmap in Figure 3, where darker green indicates higher performance. Models perform worse on benchmarks such as MedCalc-Bench (calculating medical values from patient notes), EHRSQL (generating SQL queries from natural language instructions for clinical research, originally intended as a code generation dataset), and MIMIC-IV Billing Code (assigning ICD-10 codes to clinical notes). The best performance is seen for the NoteExtract benchmark (extracting specific information from clinical notes). Broader category-level trends are described below.
Figure 3: Heatmap of normalized scores (0–1) for each model (rows) across 35 benchmarks (columns). Scores are normalized for visualization purposes; the official leaderboard reports the original (unnormalized) scores. Dark green indicates high performance; dark red indicates low performance. The statistical significance of model differences varies by benchmark, with an analysis of minimum detectable effects shown in Appendix E. Metrics include EM as exact match, Jury Score as the average normalized score from three frontier LLMs, MedCalc Accuracy as exact match or thresholded match depending on question type, MedFlagAcc as binary accuracy for detecting presence of errors, EHRSQLExeAcc as execution accuracy of generated code against a target output, and MIMICBillingF1 as the F1 score on ICD-10 codes extracted from a medical note.
@@
Figure 4 presents performance scores by the five top-level categories in the taxonomy. Most models achieve their highest mean scores in Clinical Note Generation (0.74-0.85) and Patient Communication & Education (0.76-0.89). They exhibit moderate performance in Medical Research Assistance (0.65-0.75) and Clinical Decision Support (0.61-0.76) and obtain generally lower scores in Administration & Workflow (0.53-0.63). Performance variations across categories reflect distinct task-based challenges: free-text generation tasks (Clinical Note Generation, Patient Communication) leverage models’ natural language strengths, while structured reasoning tasks (Clinical Decision Support, Administration) require domain-specific knowledge integration and logical inference. These patterns have important implications for selective deployment strategies in healthcare settings.
At the model level, DeepSeek R1 and o3-mini lead across most categories, with DeepSeek R1 excelling in Clinical Note Generation (0.85, tied with o3-mini) and Patient Communication & Education (0.89), while o3-mini leads in Clinical Decision Support (0.76) and Medical Research Assistance (0.75). The Claude Sonnet series (3.5 and 3.7) demonstrates consistent performance with scores of 0.82-0.83 in Clinical Note Generation and 0.83-0.84 in Patient Communication & Education. The GPT series (4o and 4o-mini) shows moderate consistency with scores of 0.79-0.80 in Clinical Note Generation and 0.81-0.82 in Patient Communication & Education. The Gemini models exhibit greater variation, with scores of 0.74-0.78 in Clinical Note Generation and 0.76-0.81 in Patient Communication & Education, where Gemini 2.0 Flash substantially outperforms 1.5 Pro. The open-source Llama 3.3 performs well in Patient Communication & Education (0.81) but shows the lowest score in Administration & Workflow (0.53), indicating areas for future improvement.
Figure 4: Mean normalized scores (0-1 scale) across the five categories for all evaluated models. Darker green represents higher scores. Models are ordered by mean win rate from top (highest) to bottom (lowest), while categories are arranged left to right.
@@ 2.3.3 Cost–Performance Analysis
We estimated the cost (USD) of evaluating each model based on publicly listed pricing as of 05/12/2025, using the total input tokens and maximum output tokens consumed during benchmark runs and LLM-jury evaluation. These costs are an upper-bound estimate, since models may generate fewer tokens than the maximum allowed output. We plotted the mean win-rate against the cost (Figure 5, with a detailed breakdown in Table 3, and inference costs in Appendix G). As expected, non-reasoning models—GPT-4o mini (\$805) and Gemini 2.0 Flash (\$815)—incurred the lowest costs and achieved win-rates of 0.39 and 0.42, respectively. Open-source Llama 3.3 Instruct (\$940) had a 0.30 win-rate, while Gemini 1.5 Pro (\$1,130) reached 0.24. Reasoning models incurred higher costs, DeepSeek R1 (\$1,806) and o3-mini (\$1,722), with win-rates of 0.66 and 0.64, respectively. Claude 3.5 Sonnet (\$1,571) and Claude 3.7 Sonnet (\$1,537) provide a good cost-performance balance, achieving $\sim 0.63$ win-rate at reduced costs.
Figure 5: Scatter plot of mean win-rate (y-axis) versus estimated computational cost (x-axis) for each of the nine models across 35 benchmarks. Each point represents a model, with the position indicating the relationship between performance (y-axis) and total cost of evaluation, including benchmark runs and evaluation by LLM-jury (x-axis). Costs represent upper-bound estimates based on maximum output token usage.
@@
Table 3: Comparison of LLMs by architecture, access type, and token usage metrics across MedHELM. Benchmark tokens denote total input/output tokens in completing the medical task represented by the benchmark; jury tokens are tokens used for open-ended evaluations via the LLM-jury. The total cost reflects the estimated per-model expense of running a MedHELM evaluation across 35 benchmarks and represents an upper bound based on maximum output token usage. For nine models and 35 benchmarks in the current MedHELM suite, one update to the leaderboard is estimated to cost $11,739.07$.